OpenStudy (anonymous):

What is the fourth term of (a - sqrt(2))^8

OpenStudy (tkhunny):

Mechanical hint. a 8 \(\sqrt{2}\) 0 a 7 \(\sqrt{2}\) 1 a 6 \(\sqrt{2}\) 2 a 5 \(\sqrt{2}\) 3

OpenStudy (anonymous):

i guess i dont know what that means ..

OpenStudy (tkhunny):

You probably have a fancy formula for each specific term. You should use that. \({8\choose{n}}x^{n}y^{8-n} = (x + y)^{8}\) That may be convenient to find the 133rd term from a 3645-term expansion. None terms? Just write them all out. \({8\choose{0}}\) a 8 \(\sqrt{2}\) 0 \({8\choose{1}}\) a 7 \(\sqrt{2}\) 1 \({8\choose{2}}\) a 6 \(\sqrt{2}\) 2 \({8\choose{3}}\) a 5 \(\sqrt{2}\) 3 Almost done!

OpenStudy (tkhunny):

* nine

OpenStudy (tkhunny):

(8 n )x n y 8−n =(x+y) 8 Of course, I meant \((x+y)^{8} = \sum_{n=0}^{8} {{8}\choose{n}}x^{8-n}y^{n}\) Bad typing!